What is the temperature coefficient of resistance of a piece of wire having a resistance of 11.56 ohms at 25°C and 10.22 ohms at 0°C? What will be its resistance at 50°C?


1
Expert's answer
2021-05-13T09:21:45-0400

Using the two known values of resistance, we can determine the temperature coefficient of resistance for this material:


"R_{25}=R_0(1+\alpha(T_{25}-T_0)),\\\space\\
\alpha=\frac{1}{T_{25}-T_0}\bigg(\frac{R_{25}}{R_0}-1\bigg)=5.24·10^{-3}/°\text C."

The resistance at 50°C:


"R_{50}=10.22(1+5.24·10^{-3}(50-0))=12.90\space\Omega."